OpenStudy (anonymous):

(p^2 - q^2)/(p+q) times (p-q)/(p+q) does this equal p^2-2pq+q^2?

OpenStudy (anonymous):

p^2-q^2 = (p+q)(p-q) so-> (p^2 - q^2)/(p+q) * (p-q)/(p+q) = (p-q)(p-q)/(p+q)

OpenStudy (anonymous):

= (p^2 -2pq +q^2) /(p+q)

OpenStudy (anonymous):

so would I need to simplify that into (p-q) as my final answer?

OpenStudy (anonymous):

no, it can't be simplified into (p-q) (p^2 -2pq +q^2) /(p+q) can't be simplifed because there is no common factor between the numerator and denominator

OpenStudy (anonymous):

oh ok, thanks